An automobile tire has a volume of 1000 cm³. It is inflated with carbon dioxide to pressure of 40 psig at 23°C.How many grams of carbon dioxide are contained in the tire?

1
Expert's answer
2022-01-11T12:27:02-0500

"m = \frac{P\times V\times M}{R \times T} =\frac{275.79\times 10^5 \times 0.001\times 29}{8314 \times 296} = 0.3249" Kg


 Grams of carbon dioxide is contained in the tire


M_carbon dioxide = "\frac{44}{29}\times0.3249 = 0.49 Kg"